OpenStudy (anonymous):

7.04] Choose the correct simplification of the expression the quantity a to the 3rd power over b to the 7th power all raised to the 2nd power. a to the 5th power over b to the 9th power a over b to the 14th power a to the 6th power over b to the 14th power a6b14

OpenStudy (anonymous):

\[ \frac{ a^3 }{ b^7 } ^2\]

OpenStudy (anonymous):

a to the 6th power over b to the 14th power

OpenStudy (anonymous):

when the exponents are separated by the parenthesis...you multiply...and you apply it to each term...so (a3b7)2=a3×2b7×2=a?b?

OpenStudy (anonymous):

your a ninja?

OpenStudy (anonymous):

sorry that was my brother -_-

OpenStudy (anonymous):

lol xD but i am a ninja @mjoxprm360